# concert_supp

### v2

@@ Abstract
The popularity of transfer learning stems from the fact that it can borrow information from useful auxiliary datasets. Existing statistical transfer learning methods usually adopt a global similarity measure between the source data and the target data, which may lead to inefficiency when only partial information is shared. In this paper, we propose a novel Bayesian transfer learning method named “CONCERT” to allow robust partial information transfer for high-dimensional data analysis. A conditional spike-and-slab prior is introduced in the joint distribution of target and source parameters for information transfer. By incorporating covariate-specific priors, we can characterize partial similarities and integrate source information collaboratively to improve the performance on the target. In contrast to existing work, the CONCERT is a one-step procedure, which achieves variable selection and information transfer simultaneously. We establish variable selection consistency, as well as estimation and prediction error bounds for CONCERT. Our theory demonstrates the covariate-specific benefit of transfer learning. To ensure that our algorithm is scalable, we adopt the variational Bayes framework to facilitate implementation. Extensive experiments and two real data applications showcase the validity and advantage of CONCERT over existing cutting-edge transfer learning methods.
@@ 1 Introduction
The growing emphasis on data heterogeneity in the era of big data has brought high demand for developing efficient and robust methods to integrate information from heterogeneous data sources. The homogeneous data assumption, which is commonly assumed in the traditional machine learning and statistical literature, can lead to biased inference. As a remedy, transfer learning (Pan and Yang, 2009) has emerged for its ability to borrow information from similar but different tasks, which can enhance the estimation efficiency on the target data. This is especially useful for high-dimensional data analysis, where the number of subjects in the target dataset could be much less than the number of covariates. There is a rapidly growing literature on the application of transfer learning methods in domain science, including human activity recognition (Hirooka et al., 2022), protein representation learning (Fenoy et al., 2022), and RNA modification identification (Wu et al., 2024).
One of the challenges in transfer learning lies in how to extract useful information from source datasets, and statistical research on transfer learning has grown rapidly in recent years. Among them, Li et al. (2022), Tian and Feng (2023) and Li et al. (2023) focused on (generalized) linear models, Li et al. (2023) investigated Gaussian graphical models, Zhang and Zhu (2022) studied quantile models, and Lin and Reimherr (2022) explored functional data. The notion of these methods is to measure the similarity between the source and target based on the global difference between parameters and adaptively choose informative sources with small global differences. However, the global similarity measure may lead to an all-in-or-all-out principle where each source dataset is either fully utilized for information transfer or completely disregarded, resulting in inefficiency.
To see when global similarity may fail, we illustrate two cases of partial similarity structure in the left panel of Figure 1. Figure 1(a) represents the case where sources contain heterogeneous information and only share part of the parameters with the target, and Figure 1(c) demonstrates the case where sources contain all significant signals of the target but have redundant information on additional dimensions.
Figure 1: Illustration of the two cases of partial information similarity. In (a) and (c), the true variable identities are clarified with Target Positive, Source Positive, and Negative representing influential variables in the target, transferable variables in sources, and negative variables, respectively. In (b) and (d), the corresponding estimated posterior probabilities of all variables are shown in terms of significance in the target and transferability in sources.
@@
Another challenge arises from the unknown similarities between the target and source datasets. To avoid negative transfer (Zhang et al., 2023), a two-stage framework has been developed in the transfer learning literature (Li et al., 2022; Tian and Feng, 2023; Li et al., 2023), where the global similarity between sources and the target is evaluated first. Informative sources are then selected based on case-specific criteria, or a model averaging procedure is implemented. Since the underlying similarity is not known in advance, selection methods usually rely on a global similarity gap assumption between informative and non-informative sources (Tian and Feng, 2023; Zhang and Zhu, 2022) to ensure consistent source selection. In the absence of global similarity, detecting similarity becomes more challenging. Therefore, a partially adaptive and robust transfer procedure is needed.
@@
Motivated by the above challenges, we propose a novel Bayesian transfer learning method designed to enable robust partial information transfer for high-dimensional generalized linear regression. We introduce a new conditional spike-and-slab prior in conjunction with the traditional spike-and-slab prior (George and McCulloch, 1993; Ishwaran and Rao, 2005; Ročková and George, 2018) for the joint distribution of parameters across both the target and sources. To characterize the partial similarity, we equip the coefficient of each covariate from each source with a conditional spike-and-slab prior centered at the corresponding target parameter. With this joint prior, our method can adaptively achieve variable selection and information transfer in a one-step manner. We name our proposed method CONCERT for CONditional spike-and-slab for Covariate-Elaborated Robust Transfer, which robustly transfers information from informative covariates across different sources, ensuring they can be collaboratively integrated to improve the target. In Figure 1(b) and 1(d), we show the validity of CONCERT on joint variable and similarity selection. Details can be found in Section LABEL:append:simu1 of Supplementary Material.
@@
Several independent lines of work provide reasons for Bayesian transfer learning. Among them the target prior is assigned conditionally based on source knowledge, either directly on source data (Karbalayghareh et al., 2018) or on source predictive density (Papež and Quinn, 2022; Jirsa et al., 2021), and the conditional posterior of target response is under concern. Our method differs significantly from these approaches by introducing a joint association of the parameters rather than directly assuming a joint distribution on the target and source data. The proposed method can significantly reduce model complexity while enhancing both efficiency and interpretability.
@@
The CONCERT contributes to the literature in four aspects. Firstly, we allow partial information transfer. By characterizing the covariate-specific similarity for each source using the conditional spike-and-slab prior, we can fully utilize the partially useful information from different sources even in the absence of global similarity. We showcase the advantage over existing transfer learning methods through extensive numerical experiments and real data analysis. Secondly, we establish the variable selection consistency. Specifically, we relax the minimal signal strength condition in the traditional high-dimensional variable selection literature in a covariate-specific manner. Our theory demonstrates that the more similar sources with the target on a particular variable, the lower the minimum signal strength required to detect that variable in the target. In other words, we can detect weaker signals in the target with the benefit of borrowing partial information from more sources. The estimation and prediction errors of the CONCERT also enjoy faster convergence rates than single-task Bayesian regression.
@@
Thirdly, the adoption of spike-and-slab prior for transfer learning here also provides a new perspective beyond the scope of conventional spike-and-slab prior, which is initially designed for variable selection. Under a mild local similarity gap condition, we also establish the consistency of source similarity detection under the conditional spike-and-slab prior. Lastly, we utilize variational Bayes to make our algorithm scalable with an ignorable loss of accuracy. Beyond the common mean-field variational family, a novel hierarchical structure is adopted to incorporate parameter dependence.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the problem setup and propose our CONCERT method for partial information transfer. A scalable variational Bayes algorithm is presented in Section 3. Theoretical guarantees for our method are provided in Section 4. Extensive experiments and two real data applications are provided in Sections 5 and 6, respectively. We conclude this paper in Section 7.
@@ 2 Methodology
In this section, we will introduce our CONCERT method. Consider the high-dimensional generalized linear models with response $y\in\mathbb{R}$ and covariates $x\in\mathbb{R}^{p}$. The conditional distribution of $y$ given $x$ follows
@@
for $k=1,\ldots,K$, respectively. Although additional information is provided by the samples of these source datasets, the heterogeneity between the source and the target may hinder the performance of existing methods based solely on the target data. Our aim is to extract as much useful information as possible from the source data to help improve the performance on the target while preventing negative transfer.
@@ 2.1 Spike-and-Slab for Variable Selection
Before we introduce our tailored transfer learning framework, we first provide more background on spike-and-slab for variable selection. In the Bayesian framework, the spike-and-slab prior has been widely adopted to perform variable selection, where a mixture of distributions is provided as the prior for the interested parameters. A crucial advantage of spike-and-slab is its availability of quantifying uncertainty for both parameter estimation and model selection (Narisetty et al., 2019; Nie and Ročková, 2023). Various mixture strategies have been adopted, including the absolute continuous spike-and-slab (Biswas et al., 2022) and the point-mass spike (Zeng et al., 2022). In this paper, we choose the Gaussian distribution as the slab prior and the Dirac mass at zero as the spike prior. If we only had the target data, let the priors on $\boldsymbol{\beta}^{(0)}$ be
@@
where $Z_{j}$’s are the latent variable selection indicators showing whether the $j\text{th}$ covariate of $x^{(0)}$ is informative, i.e., whether its corresponding coefficient $\beta^{(0)}_{j}$ is away from zero for $j=1,\ldots,p$, $\delta_{x}$ denotes the Dirac mass at $x$, and $\eta$ and $q_{0}$ are the hyperparameters. Based on the likelihood and priors, one can make inferences from the posterior distribution on both variable selection and parameter estimation.
@@ 2.2 Conditional Spike-and-Slab for Similarity Selection
Now we illustrate how spike-and-slab priors can also be used to transfer information from sources to the target in a covariate-specific manner. A source data is defined to be helpful for estimating the target model in the sense that the information carried by a subset of covariates is similar to that in the target data. One simple example is $\boldsymbol{\beta}^{(k)}\approx\boldsymbol{\beta}^{(0)}$ where the entire set of covariates is helpful and can be transferred. However, a more reasonable case is that only a fraction of the covariates are useful, that is, $\boldsymbol{\beta}^{(k)}_{S}\approx\boldsymbol{\beta}^{(0)}_{S}$ for some $S\subseteq\{1,\ldots,p\}$ and other covariates are not informative. Though such source data is not fully useful, we can still selectively exploit relevant information to improve target learning.
@@
$\beta^{(k)}_{j}\mid\beta^{(0)}_{j},I^{(k)}_{j}\stackrel{{\scriptstyle ind}}{{\sim}}I^{(k)}_{j}N(\beta^{(0)}_{j},\tau_{k}^{2})+(1-I^{(k)}_{j})\delta_{\beta^{(0)}_{j}},\quad I^{(k)}_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\operatorname{Bernoulli}(q_{k}),$
where $I^{(k)}_{j}$’s are the latent covariate-specific similarity indicators showing whether the $j\text{th}$ covariate of source $k$ is transferable for $j=1,\ldots,p$ and $k=1,\ldots,K$, and $\tau_{k}$ and $q_{k}$ are the hyperparameters. Our idea differs from the conventional spike-and-slab prior designed for variable selection. To better interpret our conditional spike-and-slab prior, we can define $\delta^{(k)}_{j}=\beta^{(k)}_{j}-\beta^{(0)}_{j}$ and rewrite the prior as
$\delta^{(k)}_{j}\mid\beta^{(0)}_{j},I^{(k)}_{j}\stackrel{{\scriptstyle ind}}{{\sim}}I^{(k)}_{j}N(0,\tau_{k}^{2})+(1-I^{(k)}_{j})\delta_{0},$
which inherits the spirit of mixture modeling in spike-and-slab. Notably different from the traditional spike-and-slab, the conditional spike-and-slab is designed for similarity selection, and the indicators take the role of selecting informative covariates from different sources. If $I^{(k)}_{j}=0$, the prior of $\beta^{(k)}_{j}$ concentrates on the point mass at $\beta^{(0)}_{j}$, showing a prior belief on $\beta^{(k)}_{j}=\beta^{(0)}_{j}$ and thus the corresponding covariate being transferable. Otherwise, if $I^{(k)}_{j}=1$, the prior of $\beta^{(k)}_{j}$ is diffusive away from $\beta^{(0)}_{j}$ under a large variance $\tau^{2}_{k}$, and the corresponding covariate is not considered useful for the target data.
Through the covariate-specific indicator, we can explore the partial similarity between the source and the target. Existing transfer learning methods (Li et al., 2022; Lin and Reimherr, 2022; Tian and Feng, 2023; Li et al., 2023) usually define informative sources based on the norm of difference $\boldsymbol{\beta}^{(k)}-\boldsymbol{\beta}^{(0)}$. This global similarity measure may fail to fully utilize useful information from partially similar sources, as illustrated in Figure 1. Here we adopt a new partial similarity measure to fill this gap. As we will see in our numerical studies in Section 5, our method outperforms existing methods in both the globally informative setting and the partially informative setting.
@@
In contrast to previous works in Bayesian transfer learning (Karbalayghareh et al., 2018; Papež and Quinn, 2022), our method does not impose additional assumptions on the target prior. Our target prior does not involve source information and thus remains in the original form of the sole target model, which blesses it with model consistency and interpretability despite the existence of sources.

### v3

# Covariate-Elaborated Robust Partial Information Transfer with Conditional Spike-and-Slab Prior

@@ Abstract
The popularity of transfer learning stems from the fact that it can borrow information from useful auxiliary datasets. Existing statistical transfer learning methods usually adopt a global similarity measure between the source data and the target data, which may lead to inefficiency when only partial information is shared. In this paper, we propose a novel Bayesian transfer learning method named “CONCERT” to allow robust partial information transfer for high-dimensional data analysis. A conditional spike-and-slab prior is introduced in the joint distribution of target and source parameters for information transfer. By incorporating covariate-specific priors, we can characterize partial similarities and integrate source information collaboratively to improve the performance on the target. In contrast to existing work, the CONCERT is a one-step procedure which achieves variable selection and information transfer simultaneously. We establish variable selection consistency, as well as estimation and prediction error bounds for CONCERT. Our theory demonstrates the covariate-specific benefit of transfer learning. To ensure the scalability of the algorithm, we adopt the variational Bayes framework to facilitate implementation. Extensive experiments and two real data applications showcase the validity and advantages of CONCERT over existing cutting-edge transfer learning methods.
@@ 1 Introduction
The growing emphasis on data heterogeneity in the era of big data has brought high demand for developing efficient and robust methods to integrate information from heterogeneous data sources. The homogeneous data assumption commonly assumed in the traditional machine learning and statistical literature can lead to biased inference. Alternatively, transfer learning (pan2009survey) has emerged as a powerful tool to borrow information from similar but different tasks, which can enhance the estimation efficiency on the target data. This is especially useful for high-dimensional data analysis, where the number of subjects in the target dataset could be much smaller than the number of covariates. There is a rapidly growing literature on the application of transfer learning methods in domain science, including human activity recognition (hirooka2022ensembled), protein representation learning (fenoy2022transfer), and RNA modification identification (wu2024transfer).
One of the challenges in transfer learning lies in how to extract useful information from source datasets, and statistical research on transfer learning has grown rapidly in recent years. Among them, li2022trans, tian2022transfer and li2024estimation focused on (generalized) linear models, li2023transfer investigated Gaussian graphical models, zhang2022transfer studied quantile models, and lin2022transfer explored functional data. The notion of these methods is to measure the similarity between the source and target based on the global difference between parameters and to choose informative sources with small global differences adaptively. However, the global similarity measure may lead to an all-in-or-all-out principle where each source dataset is either fully utilized for information transfer or completely disregarded, resulting in inefficiency.
To understand when global similarity may fail, we illustrate two cases of partial similarity structure in the left panel of Figure 1. Figure 1(a) represents the case where sources contain heterogeneous information and only share part of the parameters with the target, and Figure 1(c) demonstrates the case where sources contain all significant signals of the target but have redundant information on additional dimensions.
Figure 1: Illustration of the two cases of partial information similarity. In (a) and (c), the true variable identities are clarified with Target Positive, Source Positive, Target Negative, and Source Negative representing influential variables in the target, transferable variables in sources, non-influential variables in the target, and non-transferable variables in sources, respectively. In (b) and (d), the corresponding estimated posterior probabilities of all variables are shown in terms of significance in the target and transferability in sources.
@@
Another challenge arises from the unknown similarities between the target and source datasets. To avoid negative transfer (zhang2020survey), a two-stage framework has been developed in the transfer learning literature (li2022trans; tian2022transfer; li2024estimation), where the global similarity between sources and the target is evaluated first. Informative sources are then selected based on case-specific criteria, or a model averaging procedure is implemented. Since the underlying similarity is not known in advance, selection methods usually rely on a global similarity gap assumption between informative and non-informative sources (tian2022transfer; zhang2022transfer) to ensure consistent source selection. In the absence of global similarity, detecting similarity becomes more challenging. Therefore, a partially adaptive and robust transfer procedure is needed.
@@
Motivated by the above challenges, we propose a novel Bayesian transfer learning method designed to enable robust partial information transfer for high-dimensional generalized linear regression. We introduce a new conditional spike-and-slab prior in conjunction with the traditional spike-and-slab prior (george1993variable; Ishwaran2005ss; Rockova2018ss) for the joint distribution of parameters across both the target and sources. To characterize the partial similarity, we equip the coefficient of each covariate from each source with a conditional spike-and-slab prior centered at the corresponding target parameter. The spike-and-slab prior can be viewed as a Bayesian approximation to the $L_{0}$ penalty (narisetty2014bayesian), which enables partial information transfer by preserving large discrepancies between source and target domains. This avoids the over-shrinkage issue associated with the commonly used $L_{1}$ penalty, which may lead to negative transfer. We name our proposed method CONCERT for CONditional spike-and-slab for Covariate-Elaborated Robust Transfer, which transfers information from informative covariates across different sources, ensuring robust integration to improve the target. In Figure 1(b) and 1(d), we show the validity of the proposed method on joint variable and similarity selection. Details can be found in Section D.7.1 of the Supplementary Material.
@@
Several independent lines of work provide reasons for Bayesian transfer learning. Among them, the prior distribution for the target model is assigned conditionally based on source knowledge, either directly on source data (karbalayghareh2018optimal) or on source predictive density (jirsa2021bayesian; papevz2022transferring), and the conditional posterior of target response is under concern. Our method differs from these approaches by introducing a joint association of the parameters rather than directly assuming a joint distribution on the target and source data. The proposed method can significantly reduce model complexity while enhancing both efficiency and interpretability.
@@
The CONCERT brings the following contributions to the transfer learning literature. Firstly, we allow partial information transfer. By characterizing the covariate-specific similarity for each source using the conditional spike-and-slab prior, we can fully utilize the partially useful information from different sources even in the absence of global similarity. We showcase the advantage over existing transfer learning methods through extensive numerical experiments and real data analysis. Secondly, we establish the variable selection consistency. Specifically, we relax the minimal signal strength condition in the traditional high-dimensional variable selection literature in a covariate-specific manner. Our theory demonstrates that the more similar the sources are to the target on a particular variable, the lower the minimum signal strength required to detect that variable in the target. In other words, we can detect weaker signals in the target with the benefit of borrowing partial information from more sources. The estimation and prediction errors of the CONCERT also enjoy faster convergence rates than a single-task Bayesian regression.
@@
Thirdly, the adoption of spike-and-slab prior for transfer learning provides a new perspective beyond the scope of conventional spike-and-slab prior, which is initially designed for variable selection. Under a mild local similarity gap condition, we establish the consistency of source similarity detection under the conditional spike-and-slab prior. Lastly, we utilize variational Bayes to achieve scalability with a negligible loss of accuracy. Beyond the common mean-field variational family, a novel hierarchical structure is proposed to incorporate parameter dependence. Our theoretical analysis demonstrates that the VB posterior attains near-optimal contraction rates, comparable to those of the true posterior.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the problem setup and propose the CONCERT method for partial information transfer. A scalable variational Bayes algorithm is presented in Section 3. Theoretical guarantees for our method are provided in Section 4. Extensive experiments and two real data applications are provided in Sections 5 and 6, respectively. We conclude this paper in Section 7.
@@ 2 Methodology
In this section, we introduce our CONCERT method. Consider the high-dimensional generalized linear models with response $y\in\mathbb{R}$ and covariates $x\in\mathbb{R}^{p}$. The conditional distribution of $y$ given $x$ follows
@@
for $k=1,\ldots,K$, respectively. Although additional information is available from the samples of these source datasets, the heterogeneity between the sources and the target may hinder the performance of single-task methods based solely on the target data. Our aim is to extract as much useful information as possible from the source data to help improve the performance on the target while preventing negative transfer.
@@ 2.1 Spike-and-Slab for Variable Selection
Before we introduce our tailored transfer learning framework, we first provide more background on the spike-and-slab prior. In the Bayesian framework, the spike-and-slab prior has been widely adopted to perform variable selection, where a mixture of distributions is specified as the prior for the parameters of interest. Its selection properties closely resemble those of the $L_{0}$ penalty (narisetty2014bayesian), making it highly effective for variable selection in high-dimensional settings. A crucial advantage of spike-and-slab is its availability to quantify uncertainty for both parameter estimation and model selection (Narisetty2019ss; Nie2022ss). Various mixture strategies have been adopted, including the absolute continuous spike-and-slab (Biswas2022ScalableS) and the point mass spike (Zeng2022ss). In this paper, we choose the Gaussian distribution as the slab prior and the Dirac mass at zero as the spike prior. If we only had the target data, for each component $\beta^{(0)}_{j}$ of $\boldsymbol{\beta}^{(0)}$ for $j=1,\ldots,p$, let the prior be,
@@
where $Z_{j}$’s are latent variable selection indicators showing whether the $j\text{th}$ covariate of $x^{(0)}$ is informative, i.e., whether its corresponding coefficient $\beta^{(0)}_{j}$ is away from zero, $\delta_{x}$ denotes the Dirac mass at $x$, and $\eta$ and $q_{0}$ are hyperparameters. Let $S=\{j\in[p]:Z_{j}=1\}$ be the index set of the target signals. Based on the likelihood and priors, one can make inferences from the posterior distribution on both variable selection and parameter estimation.
Remark 1.
We choose point mass spike in our prior as it is preferred for variational Bayes implementation (Carbonetto2012vb; ray2022vb). A continuous spike-and-slab prior can also be considered with the implementation of MCMC approaches.
@@ 2.2 Conditional Spike-and-Slab for Similarity Selection
Now we illustrate how spike-and-slab priors can be leveraged for covariate-specific information transfer from sources to the target. A source model is considered helpful for target learning in the sense that the information carried by the covariates is similar to that in the target. A simple example is when $\boldsymbol{\beta}^{(k)}\approx\boldsymbol{\beta}^{(0)}$, meaning that the entire set of covariates is helpful and fully transferable. However, as shown in Figure 1, a more realistic scenario is that only a fraction of the covariates are useful, that is, $\boldsymbol{\beta}^{(k)}_{G}\approx\boldsymbol{\beta}^{(0)}_{G}$ for some subset $G\subseteq\{1,\ldots,p\}$, where $\boldsymbol{\beta}^{(k)}_{G}$ denotes the subvector of coefficients indexed by $G$, while the remaining covariates are not informative. Though such source data is not fully useful, we can still selectively exploit relevant information to improve target learning.
@@
$\beta^{(k)}_{j}\mid(\beta^{(0)}_{j},I^{(k)}_{j})\stackrel{{\scriptstyle ind}}{{\sim}}I^{(k)}_{j}N(\beta^{(0)}_{j},\tau_{k}^{2})+(1-I^{(k)}_{j})\delta_{\beta^{(0)}_{j}},\quad I^{(k)}_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\operatorname{Bernoulli}(q_{k}),$
where $I^{(k)}_{j}$’s are the latent covariate-specific similarity indicators showing whether the $j\text{th}$ covariate of source $k$ is transferable for $j=1,\ldots,p$ and $k=1,\ldots,K$, and $\tau_{k}$ and $q_{k}$ are the hyperparameters. To better interpret our conditional spike-and-slab prior, we can define $\boldsymbol{\delta}^{(k)}=\boldsymbol{\beta}^{(k)}-\boldsymbol{\beta}^{(0)}$, and for each component $\delta^{(k)}_{j}$ of $\boldsymbol{\delta}^{(k)}$, rewrite the prior as
$\delta^{(k)}_{j}\mid(\beta^{(0)}_{j},I^{(k)}_{j})\stackrel{{\scriptstyle ind}}{{\sim}}I^{(k)}_{j}N(0,\tau_{k}^{2})+(1-I^{(k)}_{j})\delta_{0},$
which inherits the spirit of mixture modeling in spike-and-slab. Notably different from the traditional spike-and-slab, the conditional spike-and-slab is designed for similarity selection, and the indicators take the role of selecting informative covariates from different sources. If $I^{(k)}_{j}=0$, the prior of $\beta^{(k)}_{j}$ concentrates on the point mass at $\beta^{(0)}_{j}$, showing a prior belief in $\beta^{(k)}_{j}=\beta^{(0)}_{j}$, and thus the corresponding covariate being transferable. Otherwise, if $I^{(k)}_{j}=1$, the prior of $\beta^{(k)}_{j}$ is away from $\beta^{(0)}_{j}$ with a large variance $\tau^{2}_{k}$, and the corresponding covariate is not considered useful for the target data. We denote $T_{k}=\{j\in[p]:I_{j}^{(k)}=1\}$ to be the index set of non-transferable signals for source $k$, and $T=\{T_{k}\}_{k=1}^{K}$. The covariate-specific indicators enable us to explore partial similarity between sources and the target, addressing the limitation of existing transfer learning methods (li2022trans; tian2022transfer; li2024estimation) that rely on global similarity measures such as $\|\boldsymbol{\beta}^{(k)}-\boldsymbol{\beta}^{(0)}\|_{1}$ and may overlook partially relevant sources.
Given that the spike-and-slab prior provides an approximation to the $L_{0}$ penalty, our method can be seen as replacing the $L_{1}$ penalty on $\boldsymbol{\delta}^{(k)}$ in the previous transfer learning methods with an $L_{0}$ penalty, which allows a shift from global information transfer to partial information transfer. Specifically, if certain components of $\boldsymbol{\delta}^{(k)}$ in source $k$ are large, indicating substantial differences between source $k$ and the target in certain dimensions, the $L_{1}$ penalty tends to overly suppress these differences, potentially neglecting the heterogeneity and leading to negative transfer. As a result, existing transfer learning methods often exclude such source datasets from the transferable set in advance, limiting their contribution to target learning.
In contrast, the $L_{0}$ penalty mitigates excessive shrinkage of substantial discrepancies between source and target coefficients, preserving strong heterogeneity signals instead of compressing them toward zero. This allows meaningful differences between source and target domains to be retained, enabling selective information transfer from other relevant components within the same source dataset rather than discarding it entirely.
@@
In contrast to previous works in Bayesian transfer learning (karbalayghareh2018optimal; papevz2022transferring), our method does not impose additional assumptions on the prior distribution for the target model. Our prior for target does not involve source information and thus remains in the original form of the sole target model, which blesses it with modeling consistency and interpretability despite the existence of sources.